I need help with this Logarithmic problem in a simple to understand way, thank you. 4-2e^x=-23
Answer by ankor@dixie-net.com(15622)

4-2e^x = -23
subtract 4 from both sides
-2e^x = -23 - 4
-2e^x = -27
:
Divide both sides by -2
e^x = +13.5
;
Find the nat log of both sides
ln(e^x) = ln(13.5)
which is
x*ln(e) = ln(13.5)
:
then nat log of e is 1
x*1 = ln(13.5)
:
find the ln of 13.5
x = 2.6027
:
:
You can check this solution on a good calc; enter: 4 - 2e^2.6027 results: 23.0